Answer in 500 words with introduction, concept (body), and conclusion.

Calculate the elasticity of supply if a 15 %increase in the price of soya bean oil increases its supply from 300 to 345 units



1
Expert's answer
2021-05-16T17:39:08-0400

Price elasticity of supply measures the responsiveness to the supply of a good or service after a change in its market price. Elastic means the product is considered sensitive to price changes. Inelastic means the product is not sensitive to price movements.

"Ed = \frac{250 - 200} {P2 - 8} ×\frac{P2 + 8} {250+200} = 2,"

18P2 - 144 = P2 + 8,

17P2 = 152,

P2 = 8.94.